OpenStudy (anonymous):

if x/y = 3/4 find 5x-3y/7x+2y

OpenStudy (anonymous):

THE ANSWER IS 1/10

OpenStudy (ajprincess):

\[\frac{5x-3y }{7x+2y }\] divide by y both numerator and denominator then u will get \[\LARGE\frac{ \frac{ 5x-3y }{y } }{\frac{ 7x+2y }{ y} }\] \[=\LARGE\frac{5*\frac{x}{y}-3}{7*\frac{x}{y}+2}\] Substitute 3/4 in place of x/y. \[\LARGE\frac{5*\frac{x}{y}-3}{7*\frac{x}{y}+2}\] \[=\LARGE\frac{5*\frac{3}{4}-3}{7*\frac{3}{4}+2}\]